# Improving Automated Feedback Systems for Tutor Training in Low-Resource Scenarios through Data Augmentation

### v1

@@ VIII Conclusion
In conclusion, this study demonstrated that leveraging advanced GPT models for data augmentation significantly improves the performance of explanatory feedback systems in low-resource educational settings. Our approach effectively addressed the data scarcity issue by generating high-quality synthetic labeled responses, thereby enhancing the model’s ability to generalize across various types of praise. While the method showed robust performance, we identified areas for future improvement, including more sophisticated augmentation techniques and broader testing on diverse feedback types. Overall, our findings highlight the potential of data augmentation to reduce the dependency on large labeled datasets, making automated feedback systems more scalable and practical in diverse educational contexts.
IX Acknowledgments
This research was funded by the Richard King Mellon Foundation (Grant #10851) and the Learning Engineering Virtual Institute (https://learning-engineering-virtual-institute.org/). The opinions, findings, and conclusions expressed in this paper are those of the authors alone.
@@ A-A Details of the Fine-tuning Scheme
The fine-tuning of ChatGPT models involves several key steps to adapt the pre-trained model to specific tasks or domains. In our approach, we followed the standard process of fine-tuning ChatGPT [^2] https://platform.openai.com/docs/guides/fine-tuning. Each data point (one tutor response and its corresponding label) is structured into JSON format (see Table IX and X). The JSON format begins by presenting a Lesson Principle along with two labeled examples. The Lesson Principle provides key guidelines on how to praise students effectively, such as sincerity, specificity, immediacy, and a focus on the learning process (see details in Appendix A-B and A-C). Following this, two labeled examples are included as references to guide the model in applying the principles appropriately. In the Tutor Response field, we enter each tutor’s response that needs evaluation. The Praise Type and Content field includes labels indicating the type of praise (e.g., effort-based, outcome-based or person-based) and its content. Together, these components offer context and serve as few-shot examples, enabling fine-tuning. This process was repeated for each training data point, resulting in $k$ JSON data blocks, where $k$ is the total number of training examples. Once the $k$ JSON data blocks were created, they were compiled to form the training set, which was then submitted to the GPT fine-tuning API using the default hyperparameters and settings.
TABLE IX: Input for fine-tuning GPT 3.5 model (RQ1)
@@
TABLE X: Input for fine-tuning GPT 3.5 model (RQ2)

### v2

@@ VIII Conclusion
In conclusion, this study demonstrated that leveraging advanced GPT models for data augmentation significantly improves the performance of explanatory feedback systems in low-resource educational settings. Our approach effectively addressed the data scarcity issue by generating high-quality synthetic labeled responses, thereby enhancing the model’s ability to generalize across various types of praise. While the method showed robust performance, we identified areas for future improvement, including more sophisticated augmentation techniques and broader testing on diverse feedback types. Overall, our findings highlight the potential of data augmentation to reduce the dependency on large labeled datasets, making automated feedback systems more scalable and practical in diverse educational contexts. In addition to exploring more sophisticated augmentation techniques and broader testing across diverse feedback types, future work will also focus on identifying and mitigating potential risks and biases introduced by synthetic data. Given that generative models may inadvertently reproduce or amplify existing biases, it is critical to develop robust evaluation and fairness auditing protocols to ensure that augmented datasets support equitable and reliable automated feedback systems in educational settings.
Acknowledgments
This research was funded by the Richard King Mellon Foundation (Grant #10851) and the Learning Engineering Virtual Institute (https://learning-engineering-virtual-institute.org/). The opinions, findings, and conclusions expressed in this paper are those of the authors alone. We used ChatGPT to help revise the manuscript for grammar and fluency improvements, ensuring the text is clear and natural.
@@ A-A Details of the Fine-tuning Scheme
The fine-tuning of ChatGPT models involves several key steps to adapt the pre-trained model to specific tasks or domains. In our approach, we followed the standard process of fine-tuning ChatGPT [^2] https://platform.openai.com/docs/guides/fine-tuning. Each data point (one tutor response and its corresponding label) is structured into JSON format (see Table X and XI). The JSON format begins by presenting a Lesson Principle along with two labeled examples. The Lesson Principle provides key guidelines on how to praise students effectively, such as sincerity, specificity, immediacy, and a focus on the learning process (see details in Appendix A-B and A-C). Following this, two labeled examples are included as references to guide the model in applying the principles appropriately. In the Tutor Response field, we enter each tutor’s response that needs evaluation. The Praise Type and Content field includes labels indicating the type of praise (e.g., effort-based, outcome-based or person-based) and its content. Together, these components offer context and serve as few-shot examples, enabling fine-tuning. This process was repeated for each training data point, resulting in $k$ JSON data blocks, where $k$ is the total number of training examples. Once the $k$ JSON data blocks were created, they were compiled to form the training set, which was then submitted to the GPT fine-tuning API using the default hyperparameters and settings.
TABLE X: Input for fine-tuning GPT 3.5 model (RQ1)
@@
TABLE XI: Input for fine-tuning GPT 3.5 model (RQ2)